# On the origin of circular rolls in rotor-stator flow

### v1

@@ 4.3.2 Resolvent approach to boundary forcing
The connection between the linearised time integration and the resolvent approach from (12) is the Fourier transform of the linearised governing equations. Selecting the additive forcing term $\mathbf{\hat{f}}$ corresponding to the forcing protocol of (20), and then solving (13) for $\mathbf{\hat{u}}$ will therefore yield directly the Fourier transform of the response. It can be in turn compared against the Fourier transform of the probe signals extracted from linearised time integration. The same analogy has been exploited by Cerqueira & Sipp (2014) in order to validate their input/output analysis. In our case, beyond immediate validation, this comparison is useful as it draws a link between boundary forcing (although based on linear time integration rather than the nonlinear integration allowing for spectral mixing) and the inherently bulk-based resolvent approach.
@@ 4.4 Computation of the pseudospectrum
A fundamental tool in the analysis of non-normal systems is the pseudospectrum, which is a generalisation of the concept of (eigen)spectrum. Recall that the spectrum of the linearised system $\dot{{\bm{u}}}=\mathbf{L}{\bm{u}}$ is the (complex) set of eigenvalues of the associated linear operator. If a complex number $\omega$ belongs to the spectrum $\sigma$, the resolvent $(i\omega\mathbf{I}-\mathbf{L})^{-1}$ (or expressed in the present case by (14)) is undefined, whereas it is continuous and smooth as a function of $\omega$ in the immediate neighbourhood of its pole at $i\omega$. Following Trefethen & Embree (2005) the $\varepsilon$-pseudospectrum $\sigma_{\varepsilon}$ is defined as the complex set where the norm of the resolvent operator exceeds a given value $\varepsilon^{-1}$, with $\varepsilon$ a potentially small real number :
@@
It includes the point spectrum and can be seen as its generalisation. In the current work, it is mainly used as an indicator of the strong non-normality of the underlying linearized operator. Owing to the definition 21, the cut of the pseudospectrum through the real axis directly yields the gain curves plotted in figure 3. The pseudospectrum is computed in each point, analogously to the optimal gains described in section 4.1, except that $\omega$ is allowed to be complex. Note that such computations rely on the shift-and-invert algorithm itself dependent on a purely imaginary shift parameter $s$. Contours of the pseudospectrum are reported in figure 10. As expected for a strongly non-normal operator, the isocontours of $\sigma_{\varepsilon}$ do not form concentric circles around the eigenvalues yet the contours encircle more than one eigenvalue. Still, very close locally to a given eigenvalue, the isocontours found for $Re=200$ form closed loops around specific eigenvalues, as can be seen in the top right panel in figure 10. Upon increasing $Re$, the non-normality of the linearized operator increases and the background level in the pseudospectrum grows, as shown in figure 10.
Another useful information from pseudospectra is the sensitivity of an eigenvalue to arbitrary perturbations of the operator $\mathbf{L}$ (Cerqueira & Sipp, 2014; Brynjell-Rahkola et al., 2017). As shown in chapter 28 of Trefethen & Embree (2005) for a convection-diffusion operator, high levels of non-normality typically cause iterative Arnoldi methods to converge to false eigenvalues. This is easily verified by comparing the effect of different shift values $\sigma$ in the shift-inverse Arnoldi iteration. Here, as in Cerqueira & Sipp (2014), very large level of pseudospectral contours, typically above 11, prevent ARPACK from accurately converging all the eigenvalues. This algorithmic sensibility is not only a numerical convergence problem, it also highlights a physically ambiguous situation. Individual eigenvalues, when they are not robust, do not yield a specific contribution to the dynamics. In particular the associated eigenfrequencies are not resonant, and a large response to forcing is obtained even for forcing frequencies far from the eigenfrequencies (Trefethen et al., 1993). This is associated with the presence of a large hump (rather than narrow peaks) in the gain curve. In such a case, frequent in shear flows such as jets or boundary layers, one speaks of pseudo-resonance, see e.g. Garnaud et al. (2013). Non-robust eigenvalues are easily spotted as soon as their location in the complex plane depends on the shift $\sigma$ chosen by the user, see figure 10(bottom right). Conversely, the most unstable eigenvalue $\lambda_{r}=0.011,\ \lambda_{i}=1.95$ at $Re=3000$ in figure 10 can be trusted, and the growth rate $\lambda_{r}$ predicted by the computation of eigenvalues will correspond to the growth rate of instability upon time linearized integration.
@@ 5 Experimental comparison
Numerical studies do not report sustained circular rolls in the absence of external forcing, at the exception of Daube & Le Quéré (2002) who considered larger $Re$ values.
@@
where $\delta=0.2$ (in units of $H$) corresponds to integrating over 4/5 of the interdisc spacing (thereby excluding the parallel boundary layer along the rotor). This parameter $\delta=0.2$ is preferred over $\delta=0$ for the clarity of the resulting space-time diagrams. The results for the three type of forcing are presented in figure 11. All figures are accompanied by a snapshot of the same energy $E(r,t)$ at arbitrary time, seen from above in order to emphasize the visual comparison with experimental photographs. The three types of forcing all lead to wavetrains of circular rolls with $E$ locally stronger in the interval $0.34\leq r\leq 0.6$. The rolls reaching closest to the axis correspond to monochromatic forcing, while the front seems to remain further from the axis when the spectral content of the forcing is richer. The radial velocity of the rolls is always negative (oriented towards the axis) and can be estimated directly from the space-time diagrams. Even in the monochromatic case, their velocity is not constant and it depends on $r$ : since the Bödewadt layer thickens as the perturbations migrate towards the axis where the radial velocity vanishes, the phase velocity of the rolls has to decrease, until the rolls can no longer sustain. The space-time diagram for the monochromatic case, qualitatively similar to Fig. 11 in Do et al. (2010), illustrates this point particularly well. In the harmonic case, the coexistence of different frequencies leads to different radial velocities being possible at a given radius. The dynamics becomes more complex since rolls can now travel at different velocities, leading to consecutive rolls merging or splitting. The pattern shown in Fig. 11 (middle), which shows periodic pairing and merging near $r\approx 0.75$ resembles in particular the experimental pattern shown in Figure 6 in Ref. Schouveiler et al. (2001) at the same value of $Re$. As for the white noise forcing protocol, pairing and merging events appears more intermittently admits otherwise periodic roll propagation.
@@ 6 Nonlinear receptivity
Receptivity to external disturbances in fluid flows is traditionally investigated using the linear toolbox of resolvent analysis, which is usually justified by the small perturbation velocities at play. This has the huge advantage that the responses to individual frequencies can be used as a basis to describe the global response to arbitrary forcing. Extensions of this formalism to finite-amplitude perturbations lead to new mathematical difficulties. Early efforts have focused on generalising the parabolised suitability equations (Bertolotti et al., 1992; Herbert, 1997), which unfortunately is not an option for recirculating flows. A recent attempt to add nonlinear triadic interactions to the resolvent formalism was recently proposed by Rigas et al. (2021) for boundary layer flow. For the rotor-stator flow the importance of nonlinear interactions in the dynamics of the circular rolls was already pointed out by Do et al. (2010), who measured a non-zero mean flow correction due to self-interaction of the unsteady rolls. Moreover strictly nonlinear structures (i.e. without any linear counterpart) have been determined numerically for the unforced problem by Gesla et al. (2023). It is legitimate to enquire whether they can also be detected in the forced problem, at least in the presence of large enough forcing amplitudes. In the present section, we revisit the nonlinear time integrations from Section 4 for a larger range of the parameters $A$ (forcing amplitude) and $Re$.
@@ 6.1 Effect of the nonlinearity
Figure 12: Comparison of linear and nonlinear time integration for $Re=600$, numerical resolution R1. Top: perturbation azimuthal velocity $u_{\theta}$ at location (r=5, z=0.9) from linear (left) and nonlinear time integration (right). Bottom: Corresponding spectrum of non-linear (solid lines) and linear (dashed lines) at $z=0.9$ for various $r$.
Figure 13: Comparison of linear and nonlinear time integration for $Re=2100$, numerical resolution R1. Top: perturbation azimuthal velocity $u_{\theta}$ at location (r=5, z=0.9) from linear (left) and nonlinear time integration (right). Bottom: Corresponding spectrum of non-linear (solid lines) and linear (dashed lines) at $z=0.9$ for various $r$.
@@
This analysis can be repeated for different values of $Re$. In figure 15 (left), the time-averaged $||\omega_{pert}||$ is plotted directly versus the forcing amplitude $A$. This representation allows for a direct assessment of the parameter values for which $||\omega_{pert}||$ scales linearly with $A$, which suggests linear receptivity. All the points lying away from the related straight lines are labelled as nonlinear in the $(A,Re)$ map of figure 15 (right). Beyond the simple linear/nonlinear labelling, we note that for the three largest values of $Re=2100,2400$ and $2700$ in figure 15 (left), the observable $||\omega_{pert}||$ saturates with increasing $A$ to an almost constant level. This suggests that the final state of the simulation is the same whatever $A$, in other words that the forced system is attracted towards a different region of the state space for at least $Re\gtrsim 2100$. A non-trivial chaotic attractor was reported, for the same parameters, for the unforced problem, by Daube & Le Quéré (2002) and Gesla et al. (2023), for $Re\gtrsim 1800$. This attractor is different from the laminar one and it has a different attraction basin in terms of initial condition. There is little doubt that the same attracting chaotic state is identified here in the presence of strong amplitude forcing (the initial condition being fixed). A closer zoom on the figure, especially in linear scale, would show that the saturated value of the observable depends weakly on $A$, which suggests that the associated attractor is gently sensitive to the unsteady forcing. When such an attractor is reached, the corresponding point $(A,Re)$ in the map of figure 15 (right) changes from red to blue. The points that are red correspond to the values of $(A,Re)$ where nonlinear effects do bend the gain curves (left figure), but no finite amplitude state is present and thus no saturation to any non-trivial state can occur. This leads to a 3-colour cartography of the parameter space $(A,Re)$ for the chosen forcing protocol parameterised by the amplitude parameter $A$. The notion of double threshold, popular in the shear flow transition community (Grossmann, 2000), appears here in both $A$ and $Re$ (rather than in initial perturbation amplitude and $Re$), for a forced rather than unforced problem. The boundary between linear and nonlinear behaviour in 15 (right) is consistent with an approximate fit of the form $A=O(Re^{-\alpha})$, with $\alpha\approx 10$. This large value of $\alpha$ is the signature of a very steep boundary. As for the boundary between red (nonlinear) and blue (non-trivial saturation), the present data does not suggest any simple fit. Besides the dataset is known to end at $Re\approx 3000$, beyond which all points are blue because the base flow is linearly unstable and the turbulent state is the only attractor left (Gesla et al., 2023). Rather than a cartography in the $(A,Re)$ plane, where $A$ is specific to a given forcing protocol, we plot in Figure 16 the time-averaged observable $\overline{||\omega_{pert}||}$ versus $Re$, the value of $A$ being treated implicitly. The data is represented using the same colour coding as the preceding figure, both in linear (left figure) and in logarithmic scales (right figure) . In both plots the data corresponding to the non-trivial top and lower branches (i.e. edge states) obtained in Gesla et al. (2023) are also superimposed for clarity (the same observable was used, see their figure 19. The two upper and lower branches from the unforced problem were reported to merge in a saddle-node bifurcation at $Re=Re_{SN}\approx 1800$. Interestingly, none of the red or blue data points, which correspond to the unforced problem, fall in the gap between the lower and upper branch. To our knowledge it is the first time that deterministic data from an unforced problem compares so favourably with data from the forced problem. The right panel of the figure shows the same data in logarithmic scale. There, for the choice of this observable, the boundary between linear and nonlinear behaviour obeys an approximate power-law scaling $\overline{||\omega_{pert}||}=O(Re^{-\beta})$, with $\beta\approx 3.88$.
@@ 6.2 Leaky chaotic attractors and finite lifetime dynamics
Departures from this picture can however be noted at lower $Re$ close (but above) $Re_{SN}$. Figure 18 shows the outcome several temporal simulations of the unforced problem for $Re=1840$, focusing on the time series of $||\omega_{pert}||(t)$. These time series demonstrate that the chaotic behaviour lasts for a finite time only, after which the observable rapidly reaches zero, indicating a return to the base flow. Some of the measured lifetimes are much longer than the duration of the return to the base flow, which justifies the label supertransient Lai & Tél (2011). The statistics of the lifetimes have been gathered over many such realisations of the deterministic problem, all initialised by different velocity fields drawn randomly. The cumulative distribution of the lifetimes $P(T>t)$, where $T$ is the lifetime, are shown in figure 19(left) for several values of $Re$ between 1800 and 1840. The distributions all look exponential for up to two decades, which suggests a memoryless process (Bottin & Chaté, 1998). The average value $\overline{T}$ of the lifetime is then plotted versus $Re$ in figure 19(right), where the $y$-axis is represented in logarithmic scale. The mean lifetime $\overline{T}$ increases fast with $Re$. Although no functional fit emerges from the data, the growth of $\overline{T}(Re)$ appears faster than exponential. The whole situation is reminiscent from the finite probabilities to relaminarise in three-dimensional turbulence in wall-bounded shear flows, notably the case pipe flow, where lifetimes were reported to increase super-exponentially with $Re$ (Hof et al., 2008). For the case of pipe flow, the academical issue of whether lifetime truly diverge at a finite value of $Re$ or stay finite was solving by remarking that another competing phenomenon, namely turbulence proliferation, would outweigh turbulence decay above some threshold in $Re$ (Avila et al., 2011) and keep turbulence alive for all times. There is no such equivalent here in axisymmetric rotor-stator flow, for which a scenario where lifetimes are always finite is a possibility. Nevertheless, we stress that this debate is essentially academical, the values of $T$ found in practice being so large that they usually exceed the observation times allowed for. In other words, the hypothesis that the invariant set underlying the existence of non-trivial dynamics in the forced problem is a deterministic attractor is only here for commodity : mathematically speaking it could possibly be a leaky attractor i.e. a chaotic saddle (Lai & Tél, 2011).
@@
Figure 19: Left: Cumulative distribution function $P(T>t)$ for the lifetime $T$ $P$ is estimated as the fraction of runs that did not relaminarise in the time interval $(0:t)$. Relaminarisation is decided whenever $||\omega_{pert}||<10^{-3}$. For each value of $Re$ around 200 simulations with different random initial perturbation were conducted. Right: Dependence of the mean lifetime $\overline{T}$ on $Re$. The different data illustrate the sensitivity of the mean lifetime w.r.t. the total number of runs. For a memoryless process the distributions in the left plot would be exactly exponential and the mean lifetime would be the reciprocal of their slope.

### v2

@@ 4.3.2 Resolvent approach to boundary forcing
The connection between the linearised time integration and the resolvent approach from (12) is the Fourier transform of the linearised governing equations. Selecting the additive forcing term $\mathbf{\hat{f}}$ corresponding to the forcing protocol of (20), and then solving (13) for $\mathbf{\hat{u}}$ will therefore yield directly the Fourier transform of the response. It can be in turn compared against the Fourier transform of the probe signals extracted from linearised time integration. The same analogy has been exploited by Cerqueira & Sipp (2014) in order to validate their input/output analysis. In our case, beyond immediate validation, this comparison is useful as it shows the quantitative consistency between boundary forcing (although based on linear time integration rather than the nonlinear integration allowing for spectral mixing) and the inherently bulk-based resolvent approach.
@@
An agreement between current results and the numerical study of Do et al. (2010) is noted when comparing figure 9 and figure 6 from Do et al. (2010). Both spectra are characterised by a broad curve centered around $\omega=2.4$. The position of the rolls around the mid radius of the cavity also agrees while comparing figure 8 and figure 5 from Do et al. (2010).
@@ 4.4 Computation of the pseudospectrum
A fundamental tool in the analysis of non-normal systems is the pseudospectrum, which is a generalisation of the concept of (eigen)spectrum. Recall that the spectrum of the linearised system $\dot{{\bm{q}}}=\mathbf{L}{\bm{q}}$ is the (complex) set of eigenvalues of the associated linear operator. If a complex number $\omega$ belongs to the spectrum $\sigma$, the resolvent $(i\omega\mathbf{I}-\mathbf{L})^{-1}$ (or expressed in the present case by (14)) is undefined, whereas it is continuous and smooth as a function of $\omega$ in the immediate neighbourhood of its pole at $i\omega$. Following Trefethen & Embree (2005) the $\varepsilon$-pseudospectrum $\sigma_{\varepsilon}$ is defined as the complex set where the norm of the resolvent operator exceeds a given value $\varepsilon^{-1}$, with $\varepsilon$ a potentially small real number :
@@
It includes the point spectrum and can be seen as its generalisation. In the current work, it is mainly used as an indicator of the strong non-normality of the underlying linearized operator. Owing to the definition 21, the cut of the pseudospectrum through the real axis directly yields the gain curves plotted in figure 3. The pseudospectrum is computed in each point, analogously to the optimal gains described in section 4.1, except that $\omega$ is allowed to be complex. Note that such computations rely on the shift-and-invert algorithm itself dependent on a shift parameter $s$. Contours of the pseudospectrum are reported in figure 10. As expected for a strongly non-normal operator, the isocontours of $\sigma_{\varepsilon}$ do not form concentric circles around the eigenvalues yet the contours encircle more than one eigenvalue. Still, very close locally to a given eigenvalue, the isocontours found for $Re=200$ form closed loops around specific eigenvalues, as can be seen in the top right panel in figure 10. Upon increasing $Re$, the non-normality of the linearized operator increases and the background level in the pseudospectrum grows, as shown in figure 10.
Another useful information from pseudospectra is the sensitivity of an eigenvalue to arbitrary perturbations of the operator $\mathbf{L}$ (Cerqueira & Sipp, 2014; Brynjell-Rahkola et al., 2017). As shown in chapter 28 of Trefethen & Embree (2005) for a convection-diffusion operator, high levels of non-normality typically cause iterative Arnoldi methods to converge to false eigenvalues. This is easily verified by comparing the effect of different shift values $s$ in the shift-inverse Arnoldi iteration. Here, as in Cerqueira & Sipp (2014), very large level of pseudospectral contours, typically above 11, prevent ARPACK from accurately converging all the eigenvalues. This algorithmic sensibility is not only a numerical convergence problem, it also highlights a physically ambiguous situation. Individual eigenvalues, when they are not robust, do not yield a specific contribution to the dynamics. In particular the associated eigenfrequencies are not resonant, and a large response to forcing is obtained even for forcing frequencies far from the eigenfrequencies (Trefethen et al., 1993). This is associated with the presence of a large hump (rather than narrow peaks) in the gain curve. In such a case, frequent in shear flows such as jets or boundary layers, one speaks of pseudo-resonance, see e.g. Garnaud et al. (2013). Non-robust eigenvalues are easily spotted as soon as their location in the complex plane depends on the shift $s$ chosen by the user, see figure 10(bottom right). Conversely, the most unstable eigenvalue $\lambda_{r}=0.011,\ \lambda_{i}=1.95$ at $Re=3000$ in figure 10 can be trusted, and the growth rate $\lambda_{r}$ predicted by the computation of eigenvalues will correspond to the growth rate of instability upon time linearized integration.
@@ 5 Experimental comparison
Numerical studies do not report sustained circular rolls in the absence of external forcing (Lopez et al., 2009; Poncet et al., 2009), at the exception of Daube & Le Quéré (2002) who considered larger $Re$ values.
@@
where $\delta=0.2$ (in units of $H$) corresponds to integrating over 4/5 of the interdisc spacing (thereby excluding the parallel boundary layer along the rotor). This parameter $\delta=0.2$ is preferred over $\delta=0$ for the clarity of the resulting space-time diagrams. The results for the three type of forcing are presented in figure 11. All figures are accompanied by a snapshot of the same energy $E(r,t)$ at arbitrary time, seen from above in order to emphasize the visual comparison with experimental photographs. The three types of forcing all lead to wavetrains of circular rolls with $E$ locally stronger in the interval $0.34\leq r\leq 0.6$ (radius non-dimensionalised with $R$ for consistency with Schouveiler et al. (2001)). The rolls reaching closest to the axis correspond to monochromatic forcing, while the front seems to remain further from the axis when the spectral content of the forcing is richer. The radial velocity of the rolls is always negative (oriented towards the axis) and can be estimated directly from the space-time diagrams. Even in the monochromatic case, their velocity is not constant and it depends on $r$ : since the Bödewadt layer thickens as the perturbations migrate towards the axis where the radial velocity vanishes, the phase velocity of the rolls has to decrease, until the rolls can no longer sustain. The space-time diagram for the monochromatic case, qualitatively similar to Fig. 11a in Do et al. (2010), illustrates this point particularly well. In the harmonic case, the coexistence of different frequencies leads to different radial velocities being possible at a given radius. The dynamics becomes more complex since rolls can now travel at different velocities, leading to consecutive rolls merging or splitting. The pattern shown in Fig. 11 (middle), which shows periodic pairing and merging near $r\approx 0.75$ resembles in particular the experimental pattern shown in Figure 6 in Ref. Schouveiler et al. (2001) at the same value of $Re$. It also shares some similarity with the dynamics reported in Fig 11b of Do et al. (2010) where the forcing is only monochromatic but harmonics in the response are generated by nonlinearity. As for the white noise forcing protocol, pairing and merging events appear more intermittently amidst otherwise periodic roll propagation.
@@ 6 Nonlinear receptivity
Receptivity to external disturbances in fluid flows is traditionally investigated using linear resolvent analysis, which is usually justified by the small perturbation velocities at play. This has the huge advantage that the responses to individual frequencies can be used as a basis to describe the global response to arbitrary forcing. Extensions of this formalism to finite-amplitude perturbations lead to new mathematical difficulties. Early efforts have focused on generalising the parabolised stability equations (Bertolotti et al., 1992; Herbert, 1997), which unfortunately is not an option for recirculating flows. A recent attempt to add nonlinear triadic interactions to the resolvent formalism was recently proposed by Rigas et al. (2021) for boundary layer flow. For the rotor-stator flow the importance of nonlinear interactions in the dynamics of the circular rolls was already pointed out by Do et al. (2010), who measured a non-zero mean flow correction due to self-interaction of the unsteady rolls. Moreover strictly nonlinear structures (i.e. without any linear counterpart) have been determined numerically for the unforced problem by Gesla et al. (2024). It is legitimate to enquire whether they can also be detected in the forced problem, at least in the presence of large enough forcing amplitudes. In the present section, we revisit the nonlinear time integrations from Section 4 for a larger range of the parameters $A$ (forcing amplitude) and $Re$.
@@ 6.1 Effect of the nonlinearity
Figure 12: Comparison of linear and nonlinear time integration for $Re=600$, numerical resolution R1. Top: perturbation azimuthal velocity $u_{\theta}$ at location (r=5, z=0.9) from linear (left) and nonlinear time integration (right). Bottom: Corresponding spectrum of non-linear (solid lines) and linear (dashed lines) velocity signal at $z=0.9$ for various $r$.
Figure 13: Comparison of linear and nonlinear time integration for $Re=2100$, numerical resolution R1. Top: perturbation azimuthal velocity $u_{\theta}$ at location (r=5, z=0.9) from linear (left) and nonlinear time integration (right). Bottom: Corresponding spectrum of non-linear (solid lines) and linear (dashed lines) velocity signal at $z=0.9$ for various $r$.
@@
This analysis can be repeated for different values of $Re$. In figure 15 (left), the time-averaged $||\omega_{pert}||$ is plotted directly versus the forcing amplitude $A$. This representation allows for a direct assessment of the parameter values for which $||\omega_{pert}||$ scales linearly with $A$, which suggests linear receptivity. All the points lying away from the related straight lines are labelled as nonlinear in the $(A,Re)$ map of figure 15 (right). Beyond the simple linear/nonlinear labelling, we note that for the three largest values of $Re=2100,2400$ and $2700$ in figure 15 (left), the observable $||\omega_{pert}||$ saturates with increasing $A$ to an almost constant level. This suggests that the final state of the simulation is the same whatever $A$, in other words that the forced system is attracted towards a different region of the state space for at least $Re\gtrsim 2100$. A non-trivial chaotic attractor was reported, for the same parameters, for the unforced problem, by Daube & Le Quéré (2002) and Gesla et al. (2024), for $Re\gtrsim 1800$. This attractor is different from the laminar one and it has a different attraction basin in terms of initial condition. There is little doubt that the same attracting chaotic state is identified here in the presence of strong amplitude forcing (the initial condition being fixed). A closer zoom on the figure, especially in linear scale, would show that the saturated value of the observable depends weakly on $A$, which suggests that the associated attractor is gently sensitive to the unsteady forcing. When such an attractor is reached, the corresponding point $(A,Re)$ in the map of figure 15 (right) changes from red to blue. The points that are red correspond to the values of $(A,Re)$ where nonlinear effects do bend the gain curves (left figure), but no finite amplitude state is present and thus no saturation to any non-trivial state can occur. This leads to a 3-colour cartography of the parameter space $(A,Re)$ for the chosen forcing protocol parameterised by the amplitude parameter $A$. The notion of double threshold, popular in the shear flow transition community (Grossmann, 2000), appears here in both $A$ and $Re$ (rather than in initial perturbation amplitude and $Re$), for a forced rather than unforced problem. The boundary between linear and nonlinear behaviour in 15 (right) is consistent with an approximate fit of the form $A=O(Re^{-\alpha})$, with $\alpha\approx 10$. This large value of $\alpha$ is the signature of a very steep boundary. As for the boundary between red (nonlinear) and blue (non-trivial saturation), the present data does not suggest any simple fit. Besides the dataset is known to end at $Re\approx 3000$, beyond which all points are blue because the base flow is linearly unstable and the turbulent state is the only attractor left (Gesla et al., 2024). Rather than a cartography in the $(A,Re)$ plane, where $A$ is specific to a given forcing protocol, we plot in Figure 16 the time-averaged observable $\overline{||\omega_{pert}||}$ versus $Re$, the value of $A$ being treated implicitly. The data is represented using the same colour coding as the preceding figure, both in linear (left figure) and in logarithmic scales (right figure). In both plots the data corresponding to the non-trivial top and lower branches (i.e. edge states) obtained in Gesla et al. (2024) are also superimposed for clarity (the same observable was used, see their figure 19). The two upper and lower branches from the unforced problem were reported to merge in a saddle-node bifurcation at $Re=Re_{SN}\approx 1800$. Interestingly, none of the red or blue data points, which correspond to the unforced problem, fall in the gap between the lower and upper branch. This suggests that the data from the deterministic problem can be used to delineate between the two types of response, even in the presence of a finite-amplitude forcing. The right panel of the figure shows the same data in logarithmic scale. There, for the choice of this observable, the boundary between linear and nonlinear behaviour obeys an approximate power-law scaling $\overline{||\omega_{pert}||}=O(Re^{-\beta})$, with $\beta\approx 3.88$.
@@ 6.2 Leaky chaotic attractors and finite lifetime dynamics
Departures from this picture can however be noted at lower $Re$ close (but above) $Re_{SN}$. Figure 18 shows the outcome several temporal simulations of the unforced problem for $Re=1840$, focusing on the time series of $||\omega_{pert}||(t)$. These time series demonstrate that the chaotic behaviour lasts for a finite time only, after which the observable rapidly reaches zero, indicating a return to the base flow. Some of the measured lifetimes are much longer than the duration of the return to the base flow, which justifies the label supertransient (Lai & Tél, 2011). The statistics of the lifetimes have been gathered over many such realisations of the deterministic problem, all initialised by different velocity fields drawn randomly. The cumulative distribution of the lifetimes $P(T>t)$, where $T$ is the lifetime, are shown in figure 19 for several values of $Re$ between 1800 and 1860. The distributions all look exponential for up to two decades, which suggests a memoryless process (Bottin & Chaté, 1998).
As noted by Avila et al. (2010), the perturbations need a certain time $t_{0}$, before they reach a leaky attractor. The mean lifetime $\tau$ of only these perturbations that have reached the leaky attractor can be computed by discarding all of the simulations that relaminarised before the time $t_{0}$:
$\tau(t_{0})=\left<t-t_{0}|t>t_{0}\right>$ (25)
[…]
In case of relaminarisation of every simulation out of $r$ total simulations, the Maximum Likelihood Estimator (MLE) of $\tau$ can be computed using
$\tau=\frac{1}{r}\sum_{i=1}^{r}t_{i}.$ (26)
where $t_{i}$ is the relaminarisation time of individual simulations. However, because of the finite numerical integration time, not all simulations relaminarise. To account for the perturbations whose lifetime exceeds a censoring time $t_{r}$, the MLE is modified as (Lawless, 2003)
$\tau=\frac{1}{r}\left(\sum_{i=1}^{r}t_{i}+(n-r)t_{r}\right),$ (27)
with $r$ thenumber of simulations that relaminarised and $n$ the total number of simulations. A 95% confidence interval of the estimator (27) is
$\tau\times\left[\frac{2r}{\chi^{2}_{2r,0.975}},\frac{2r}{\chi^{2}_{2r,0.025}}\right].$ (28)
where $\chi^{2}_{m,p}$ is the $p$th quantile of the chi-squared distribution with $m$ degrees of freedom.
The total number of simulations performed for each $Re$, together with $t_{0}$, the escape rate from the saddle $\kappa$ (Tél & Lai, 2008), its confidence interval and a corresponding $\tau$, are all reported in table 3. The dependence of $\kappa$ on $t_{0}$ and $Re$ is shown in figure 20. Linear dependence $log(\kappa(Re))$ would suggest an exponential scaling of the mean lifetime as a function of $Re$, at least over the interval in $Re$ where the data was gathered (which lies well below the linear instability threshold).
$Re$ $n$ $r$ $t_{0}$ $\kappa(t_{0})$ $\kappa$ 95% conf. int. $\tau$
1800 299 299 721.3 3.016e-03 [2.611e-03,3.449e-03] 331.6
1810 299 299 872.7 1.929e-03 [1.670e-03,2.206e-03] 518.4
1820 299 299 1026.1 1.253e-03 [1.085e-03,1.433e-03] 798.3
1830 298 293 1286.5 5.775e-04 [4.989e-04,6.618e-04] 1731.6
1840 298 261 2159.5 3.193e-04 [2.719e-04,3.704e-04] 3132.1
1850 198 114 722.6 1.169e-04 [9.640e-05,1.393e-04] 8556.8
1860 198 54 784.9 4.580e-05 [3.441e-05,5.880e-05] 21833.8
Table 3: Statistics of lifetimes. For each value of $Re$, $r$ is the number of simulations out of total $n$ to relaminarise before the censoring time $t_{r}=7800$. The distribution of lifetimes is approximately exponential for $t>t_{0}$. The maximum likelihood estimator of a mean lifetimes is computed using (27) and its confidence interval using (28). The escape rate from the saddle $\kappa=1/\tau$ and its confidence interval are plotted in figure 20.
@@
Figure 19: Cumulative distribution function $P(T>t)$ for the lifetime $T$ $P$ is estimated as the fraction of runs that did not relaminarise in the time interval $(0:t)$. Relaminarisation is decided whenever $||\omega_{pert}||<10^{-3}$. For each value of $Re$ simulations with different random initial perturbation were conducted.
[…]
The whole situation is reminiscent from the finite probabilities to relaminarise in three-dimensional turbulence in wall-bounded shear flows, notably the case pipe flow, where lifetimes were reported to increase super-exponentially with $Re$ (Hof et al., 2008). For the case of pipe flow, the issue of whether lifetime truly diverge at a finite value of $Re$ or stay finite was solved by remarking that another competing phenomenon, namely turbulence proliferation, would outweigh turbulence decay above some threshold in $Re$ (Avila et al., 2011) and keep turbulence alive for all times. There is no such equivalent here in axisymmetric rotor-stator flow, for which a scenario where lifetimes are always finite is a possibility below the threshold of linear stability of the base flow $Re_{c}\approx 3000$ (Daube & Le Quéré, 2002). Nevertheless, the values of $T$ found in practice are so large that they usually exceed the observation times allowed for. In other words, the hypothesis that the invariant set underlying the existence of non-trivial dynamics in the forced problem is a deterministic attractor is only here for commodity : mathematically speaking it could possibly be a leaky attractor i.e. a chaotic saddle (Lai & Tél, 2011).